OpenStudy (anonymous):

ken has three more coins that twice the number javier has. khald has 5 fewer coins than javier. they have 50 coins together

OpenStudy (anonymous):

how can i do this

OpenStudy (anonymous):

Let j= the number of coins Javier has, let k= the number of coins Ken has, and Let x= the number of coins khald has. Firstly, j+k+x=50. Next, 2j+3=k. Finally j=x+5. To solve a 3 equation system you must use substitution. First step: choose two equations that you have known values for. Let's chose the last two equations. Since we know j=x+5, we can substitute it into the first of the later, 2(x+5)+3=k. FInally we know k and j so we can substitute these into the first equation to find x. (x+5)+2(x+5)+3+(x)=50 Now simplify to find x. x+5+2x+10+3+x=50. 4x+18=50 4x=32 x=8 Substitute 8 back into the other equations. j=8+5; j=13. 2(13)+3=k. k=29. Khald has 8 coins, Javier has 13 coins, and Ken has 29 coins. Done :)

OpenStudy (anonymous):

thank u s0 much

OpenStudy (anonymous):

No problem would you mind rating me the best answer? :)

OpenStudy (anonymous):

where can i do that

OpenStudy (anonymous):

on my reply click best answer, top right